OpenStudy (fanduekisses):

Did I do this right? xy=4 Find dy/dt when x=8 given dx/dt=10

OpenStudy (fanduekisses):

dx/dt*y + x**dy/dt = 0 10y + 8 dy/dt = 0 8 dy/dt = -10y dy/dt = -5/4 y

OpenStudy (anonymous):

i dont think so hold on let me do it on paper

OpenStudy (perl):

dx/dt * y + x * dy/dt = 0 you did that part correctly

OpenStudy (perl):

the only thing you need is to solve for y

OpenStudy (fanduekisses):

Solve for y? why

OpenStudy (perl):

dy/dt = -5/4 y , you can plug in y, since we know that x*y = 4 , and x = 8 , that means

OpenStudy (perl):

8*y = 4 , y = 4/8

OpenStudy (fanduekisses):

ooooooooooh

OpenStudy (fanduekisses):

so dy/dt= -5?

OpenStudy (fanduekisses):

I mean -5/8